OpenStudy (anonymous):

The hubble space telescope is orbiting earth 600 km above earths surface. earths radius is about 6370 km use the pythagorean theroem to find the distance x from the telescope to earths horizon. CAn you write out the thereom and what a would be? i got a+6370=6970 all squared of course.

OpenStudy (anonymous):

what is meant by horizon ?

OpenStudy (anonymous):

is there an attached grapg or something ?

OpenStudy (anonymous):

imagince a cirlce with a triangle half way out of the cirlce. 660 km is out of the cirlce.

OpenStudy (anonymous):

sorry 600km

OpenStudy (anonymous):

so we have a right triangle , one leg = 600 +6370 = 6970 other leg is same raduis = 6370 use pyth. x^2 = 6970^2 + 6370^2 x = 9442.34

OpenStudy (anonymous):

I drew it out if that helps.

OpenStudy (anonymous):

Thanks!!!:)

OpenStudy (anonymous):

u sure what u drew is right ?

OpenStudy (anonymous):

well according to what u drew (i dont think its true) but here we go 6970^2 = 6370^2 + x^2 x = 8004

OpenStudy (anonymous):

Hi Adellable,it's a little bit difficult, one would imagine the radius represents the adjacent of the triangle, and so we are left with two sides. Now if we take the distance from earth of hubble at an angle, then it would represent the hypoteneuse, however this leaves a problem with the opposite side of the triangle, as it would be distance from hubble to centre of earth which is not what we are looking for. but if you minus the radius you would get the proper distance.

OpenStudy (anonymous):

hmmm okay thanks for your help

OpenStudy (anonymous):

\[(600+6730)^{2}=6730^{2}+x ^{2}\]

OpenStudy (anonymous):

\[7330^{2}=6730^{2}+x ^{2}\]

OpenStudy (anonymous):

\[7330^{2}-6730^{2}=x ^{2}\]

OpenStudy (anonymous):

\[53728900-45292900=x ^{2}\]

OpenStudy (anonymous):

\[8436000=x ^{2}\]

OpenStudy (anonymous):

\[x=\sqrt{8436000}\]

OpenStudy (anonymous):

x=2904.479299 This calculation is based on the diagram you supplied, where hypoteneuse= (600+6730), adjacent (or radius)=6730 and opposite=x.